Rust contains iron(111)oxide,Fe2o3 .how many moles is iron (111) ions are present in 4.0 of rust?
1
Expert's answer
2013-12-21T09:00:24-0500
Fe2O3= 2Fe(3+) + 3O2(2-)
The number of moles of iron ions is two times greater than that of iron oxide. The molecular weight M(Fe2O3)=56*2+16*3=160g/mol
n(Fe)=(4/160)*2=0.05mol